OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 2x – 2y – 2z = –8 x + 5y – 3z = 2

OpenStudy (anonymous):

2x-3x+z=-12, x-2y-2z=-8, x+5y-3z=2